The base of a rectangle is 3 1/3 feet. The area of this rectangle is 50 square ft. Marks claims the height of the rectangle is 166 2/3 feet. What is the correct height?
What is marks error?

Answer by edjones(8007)

3 1/3 =10/3
L*w=A
10x/3=50
x=50*3/10
=15'
.
Ed